In a spring block system the oscillation frequency is 8.00 Hz. If the oscillation amplitude is 0.400 m, determine the position, speed and acceleration at t = 8.00 s.
Answer #1

x = A cos ( wt)

x = 0.4* cos ( 2* 3.14* 8* 8)

x = 0.3917 m

=======

v = dx/dt = - A w sin (wt)

v = - 0.4* 2* 3.14* 8* sin ( 2 * 3.14* 8* 8)

v = - 4.068 m/s

=======

a = dv/dt = - A w^2 cos ( wt)

a = - 0.4* (2*3.4* 8)^2 cos ( 2*3.14 * 8* 8)

a = - 988.71 m/s^2

=======

Comment in case any doubt, will reply for sure.. Goodluck